OpenStudy (anonymous):

$$\LARGE \int {(x + 1){{({x^2\;+}2x + 3)}^3}\,\,dx} $$ \[\LARGE \text{ATEMPTED}\] (I followed wolframalpha's steps too ...): \[\LARGE u=x^2+2x+3\] and \[\LARGE du=u'\;dx\] \[\LARGE du=(x^2+2x+3)'\;dx\] \[\LARGE du=2x+2\;dx\] now (wolframa) when substitutes it goes like: \[\LARGE \frac12 \;\int\limits u^2 \;du\] But I'm extremely confused ... !! Please show me steps too...

OpenStudy (mimi_x3):

\[\int\limits(x+1)u^{3}*\frac{du}{2x+2} =>\int\limits(x+1)u^{3}*\frac{du}{2(x+1)} =>\frac{1}{2} \int\limits u^{3} du\]

OpenStudy (anonymous):

Thank you very much, 100% got it ;) @Mimi_x3

OpenStudy (mimi_x3):

yw (: